Science and technology. Fungicides account for 1/10 of the pesticides used in the United States. Insecticides account for 1/4 of all the pesticides used in the United States. The ratio of fungicides to insecticides used in the United States can be written 1/10 divided by 1/4. Write this ratio in simplest form.

Answer by stanbon(48558)

1/10 divided by 1/4. Write this ratio in simplest form.
--------------
(1/10)/(1/4)
Invert the denominator and multiply:
(1/10) * (4/1) = 4/10 = 2/5
=================
Cheers,
Stan H.